OpenStudy (justuu):

Medal Which measurement is most accurate to describe the amount a coffee cup could hold? 350 mL 2.4 L 25 mL 0.8 L

OpenStudy (danjs):

do you know what a 2 liter pop bottle looks like

OpenStudy (anonymous):

Lol thats not the size of a cofee cup ;)

OpenStudy (danjs):

right, you arent supposed to give out direct answers

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

I deleted it

OpenStudy (danjs):

I am trying to think of something you are familiar with to put each of these relative volumes into perspective.

OpenStudy (danjs):

2.4 L is larger than a 2 liter pop bottle, that for sure is much more than a coffee cup worth

OpenStudy (anonymous):

Well nvm she closed it...

OpenStudy (danjs):

Even 8/10 or 0.8 Liters is more than a single cup of coffee

OpenStudy (anonymous):

yep, and 25 ml is 1.5 table spoons

OpenStudy (danjs):

So it is between 350 mL and 25 mL 350 mL, is 350/1000 = 35/100 = 7/20 of a liter....that one seems the most reasonable

OpenStudy (anonymous):

It has to be 350 mL, even that is alot for a cup of coffee though